The table below gives the quantity purchased by an individual at various prices. Price Qd What is the consumer surplus for this individual if the market price is $4.50? A) 10 $10 $8 $6 $5 $4 $3 $2 B) $11 C) $11.50 D) S12 E) None of the above
Answer #1

Here the maximum willingness to pay for a customer is more than 4.5 for 4 consumers only, so only those people will buy .

The consumer surplus will be 10+8+6+5-4.5x4=29-18=11

So B will be the answer.